# Stochastic Learning of Computational Resource Usage as Graph Structured Multimarginal Schrödinger Bridge

### v1

@@ I Introduction
  • Mathematical formulations for the stochastic learning of computational resource usage as graph-structured MSBPs where the graph structures arise from single or multi-core computational platforms. These formulations are motivated by maximum likelihood interpretations in the path space.
  • Numerical algorithms for solving the proposed graph-stuctured MSBPs. We show that unlike generic graph structures, the MSBPs induced by single or multi-core computational resource usage lead to structured tensor optimization problems that can be leveraged to reduce the learning algorithms’ computational complexity from exponential to linear in number of observations.
  • Numerical experiments to illustrate the proposed learning framework. This includes a single-core MPC benchmark, and a synthetic multi-core software benchmark.
Organization
In Sec. II, we set up basic notations and background ideas to be used throughout this work. Sec. III outlines the overall problem formulation. Expanding on the same, Sec. IV explains the maximum likelihood interpretations motivating the ideas, and then gives mathematically precise formulations for single core (Sec. IV-C) and multi-core (Sec. IV-D, IV-E) CPU usages. Sec. V details the proposed algorithms and their computational complexities for solving the formulations in Sec. IV. In particular, algorithmic specifics for Sec. IV-C are in Sec. V-A, those for Sec. IV-D are in V-B, and those for Sec. IV-E are in Sec. V-C. Numerical experiments highlighting the efficacy of the proposed method are detailed in Sec. VI for both single and multi-core CPUs. Sec. VII concludes the paper. Proofs are given in the Appendix.
@@ V Algorithms
have guaranteed linear rate of convergence, and the minimizer $\bm{M}^{\rm{opt}}$ for (34) is given in terms of the converged $\bm{U}$ as
@@ V-C Projections for Series-Parallel Graph-structured MSBP
$A_{k}:=K^{k,1}\left(\prod_{\sigma=2}^{s-1}\mathrm{diag}(\bm{u}_{\sigma}^{k})K^{k,\sigma}\right),$ (52)
the projection (27) takes the form
${\rm{proj}}_{(1,1)}\!\left(\bm{K}\odot\bm{U}\right)\!$ $=\bm{u}_{1}^{1}\odot\left(\prod_{k=1}^{J}A_{k}\right)\bm{u}_{s}^{1},$ (53a)
${\rm{proj}}_{(1,s)}\!\left(\bm{K}\odot\bm{U}\right)\!$ $=\left({\bm{u}_{1}^{1}}^{\top}\prod_{k=1}^{J}A_{k}\right)^{\top}\odot\bm{u}_{s}^{1},$ (53b)
${\rm{proj}}_{(j,\sigma)}\!\left(\bm{K}\odot\bm{U}\right)\!$ $=\Bigg{(}{\bm{u}_{1}^{1}}^{\top}\bigg{(}\prod_{k=1}^{j-1}A_{k}\bigg{)}K^{j,1}$
$\bigg{(}\prod_{m=2}^{\sigma-1}\mathrm{diag}(\bm{u}_{m}^{j})K^{j,m}\bigg{)}\Bigg{)}^{\top}\odot\bm{u}_{\sigma}^{j}\odot$
$\Bigg{(}K^{j,\sigma}\bigg{(}\prod_{m=\sigma+1}^{s-1}$ $\mathrm{diag}(\bm{u}_{m}^{j})K^{j,m}\bigg{)}\bigg{(}\prod_{k=j+1}^{J}A_{k}\bigg{)}\bm{u}_{s}^{1}\Bigg{)},$ (53c)
@@
${\rm{proj}}_{(1,1),(j,2)}\!\left(\bm{K}\odot\bm{U}\right)\!=\mathrm{diag}\left(\bm{u}_{1}^{1}\right)\cdot\bigg{(}\prod_{k=1}^{j-1}A_{k}\bigg{)}K^{j,1}$
$\cdot\mathrm{diag}\left(\bm{u}_{2}^{j}\odot K^{j,2}\bigg{(}\prod_{k=3}^{s-1}\mathrm{diag}(\bm{u}_{k}^{j})K^{j,k}\bigg{)}\bigg{(}\prod_{k=j+1}^{J}A_{k}\bigg{)}\bm{u}_{s}^{1}\right),$ (54a)
${\rm{proj}}_{(j,s-1),(1,s)}\!\left(\bm{K}\odot\bm{U}\right)\!=\mathrm{diag}\Bigg{(}\Bigg{(}{\bm{u}_{1}^{1}}^{\top}\bigg{(}\prod_{k=1}^{j-1}A_{k}\bigg{)}K^{j,1}$
$\bigg{(}\prod_{k=2}^{s-2}\mathrm{diag}(\bm{u}_{k}^{j})K^{j,k}\bigg{)}\Bigg{)}^{\top}\odot\bm{u}_{s-1}^{j}\Bigg{)}\cdot K^{j,s-1}\bigg{(}\prod_{k=j+1}^{J}A_{k}\bigg{)}$
$\cdot\mathrm{diag}\left(\bm{u}_{s}^{1}\right),$ (54b)
${\rm{proj}}_{(j,\sigma),(j,\sigma+1)}\!\left(\bm{K}\odot\bm{U}\right)\!=\mathrm{diag}\Bigg{(}\Bigg{(}{\bm{u}_{1}^{1}}^{\top}\bigg{(}\prod_{k=1}^{j-1}A_{k}\bigg{)}K^{j,1}$
$\bigg{(}\prod_{k=2}^{\sigma-1}\mathrm{diag}(\bm{u}_{k}^{j})K^{j,k}\bigg{)}\Bigg{)}^{\top}\odot\bm{u}_{\sigma}^{j}\Bigg{)}\cdot K^{j,\sigma}\cdot\mathrm{diag}\Bigg{(}\bm{u}_{\sigma+1}^{j}$
$\odot K^{j,\sigma+1}\Bigg{(}\prod_{j=\sigma+2}^{s-1}\mathrm{diag}(\bm{u}_{k}^{j})K^{j,k}\bigg{)}\bigg{(}\prod_{k=j+1}^{J}A_{k}\bigg{)}\bm{u}_{s}^{1}\Bigg{)}.$ (54c)
Directly computing (52) requires $1+2(s-2)$ matrix-matrix multiplications, each in this case involving $n^{3}$ floating-point operations. However, since all unimarginal projections result in a vector, (53) involves only matrix-vector multiplications ($n^{2}$ operations) and elementwise vector multiplication ($n$ operations). We can see that (53a) and (53b) each involve
$n+J\left(1+2(s-2)\right)n^{2}$ (55)
@@
$2n+J\left(1+2(s-2)\right)n^{2}-n^{2}$ (56)
such operations.
@@
In contrast, the expression (56)$=(2Js-3J-1)n^{2}+2n\geq n^{2}+2n$. For $n$ large, both complexities are $\mathcal{O}(n^{2})$.
@@ VI-A1 Profiling
Remark 6.
In this paper, we are interested in capturing only hardware-level stochasticity, which is why in all experiments, we ensure that the cyber-physical context in which the software runs, remains fixed. For example, we can expect the NMPC to be more resource-intensive when tracking a complex reference trajectory as opposed to a simple trajectory such as a straight line. Thus, if the path to be followed by the NMPC control software is random, we can expect to capture the stochasticity of the path generation code. Similarly, if we do not fix the LLC and memory bandwidth allocated to the software, we are capturing the stochasticity of how the underlying operating system allocates resources. As we desire rather to control for these factors, we always fix both the data the software takes as initial input (the physical context) and the hardware resources available to it throughout execution (the cyber context).
@@ VI-A2 Numerical Results
We computed (57) as the square roots of the optimal values of the corresponding Kantorovich linear programs [58, Ch. 3.1] that results from specializing (29) with $s=2$, $\varepsilon=0$. As expected, placing the marginals closer together in time results in higher accuracy in predictions.
@@ VI-B2 Numerical Results
Remark 7.
For this experiment, our choice of marginals and interpolation times was made primarily for the sake of presentability – it would certainly be feasible to naïvely partition the profiles into equispaced intervals and place marginals at their boundaries, and to interpolate at will. As discussed in Sec. VI-B for the single-core experiment, placing more of these marginals temporally closer together would result in more accurate predictions.
@@
$1$ $4.079\times 10^{-5}$ $1.062\times 10^{-7}$ $2.173\times 10^{-7}$ $2.110\times 10^{-7}$ $1.575\times 10^{-7}$
$2$ 00 $1.048\times 10^{-7}$ $2.333\times 10^{-7}$ $7.303\times 10^{-8}$ $9.186\times 10^{-8}$
$3$ 00 $1.065\times 10^{-7}$ $1.576\times 10^{-7}$ $6.394\times 10^{-8}$ $6.935\times 10^{-8}$
$4$ 00 $5.553\times 10^{-8}$ $2.269\times 10^{-7}$ $1.857\times 10^{-7}$ $9.691\times 10^{-8}$
@@ VII Conclusions
We expect the proposed learning framework to be of interest to both control and real-time systems community because it opens the door for fine-grained dynamic scheduler design, especially for resource-constrained environments such as controlled cyber-physical systems. The learning-scheduling co-design that builds on the proposed method, will comprise our future work. Another direction of interest is to account for the asynchronous nature of the profiles by formulating the corresponding MSBPs with joint space-time stochasticity, i.e., as distributional generalization of optimal stopping problems initiated in the recent work [60].
@@
$={\bm{u}_{1}^{1}}^{\top}\prod_{k=1}^{J}\Bigg{(}\underbrace{K^{k,1}\left(\prod_{\sigma=2}^{s-1}\mathrm{diag}(\bm{u}_{\sigma}^{k})K^{k,\sigma}\right)}_{:=A_{k}}\Bigg{)}\bm{u}_{s}^{1}$ (63)
$={\bm{u}_{1}^{1}}^{\top}\bigg{(}\prod_{k=1}^{j-1}A_{k}\bigg{)}A_{j}\bigg{(}\prod_{k=j+1}^{J}A_{k}\bigg{)}(\bm{u}_{s}^{1})$
$=\underbrace{{\bm{u}_{1}^{1}}^{\top}\bigg{(}\prod_{k=1}^{j-1}A_{k}\bigg{)}\Bigg{(}K^{j,1}\bigg{(}\prod_{m=2}^{\sigma-1}\mathrm{diag}(\bm{u}_{m}^{j})K^{j,m}\bigg{)}}_{\bm{w}_{1}^{\top}}\cdot\mathrm{diag}(\bm{u}_{\sigma}^{j})$
$\quad\cdot\underbrace{K^{j,\sigma}\bigg{(}\prod_{m=\sigma+1}^{s-1}\mathrm{diag}(\bm{u}_{m}^{j})K^{j,m}\bigg{)}\Bigg{)}\bigg{(}\prod_{k=j+1}^{J}A_{k}\bigg{)}\bm{u}_{s}^{1}}_{\bm{w}_{2}}$ (64)
@@
$\langle\mathbf{K},\mathbf{U}\rangle=\underbrace{\mathbf{1}^{\top}}_{\bm{w}_{1}^{\top}}\cdot\mathrm{diag}(\bm{u}_{1}^{1})\cdot\underbrace{\bigg{(}\prod_{k=1}^{j-1}A_{k}\bigg{)}K^{j,1}}_{\Phi}\cdot\mathrm{diag}(\bm{u}_{2}^{j})$
$\quad\cdot\underbrace{K^{j,2}\bigg{(}\prod_{k=3}^{s-1}\mathrm{diag}(\bm{u}_{k}^{j})K^{j,k}\bigg{)}\bigg{(}\prod_{k=j+1}^{J}A_{k}\bigg{)}\bm{u}_{s}^{1}}_{\bm{w}_{3}}$ (65)
$=\underbrace{{\bm{u}_{1}^{1}}^{\top}\bigg{(}\prod_{k=1}^{j-1}A_{k}\bigg{)}K^{j,1}\bigg{(}\prod_{k=2}^{s-2}\mathrm{diag}(\bm{u}_{k}^{j})K^{j,k}\bigg{)}}_{\bm{w}_{1}^{\top}}$
$\quad\cdot\mathrm{diag}(\bm{u}_{s-1}^{j})\cdot\underbrace{K^{j,s-1}\bigg{(}\prod_{k=j+1}^{J}A_{k}\bigg{)}}_{\Phi}\cdot\mathrm{diag}(\bm{u}_{s}^{1})\cdot\underbrace{\mathbf{1}}_{\bm{w}_{3}}$ (66)
$=\underbrace{{\bm{u}_{1}^{1}}^{\top}\bigg{(}\prod_{k=1}^{j-1}A_{k}\bigg{)}K^{j,1}\bigg{(}\prod_{k=2}^{\sigma-1}\mathrm{diag}(\bm{u}_{k}^{j})K^{j,k}\bigg{)}}_{\bm{w}_{1}^{\top}}\cdot\mathrm{diag}(\bm{u}_{\sigma}^{j})$
$\quad\cdot\underbrace{K^{j,\sigma}}_{\Phi}\cdot\mathrm{diag}(\bm{u}_{\sigma+1}^{j})$
$\cdot\underbrace{K^{j,\sigma+1}\Bigg{(}\prod_{j=\sigma+2}^{s-1}\mathrm{diag}(\bm{u}_{k}^{j})K^{j,k}\bigg{)}\bigg{(}\prod_{k=j+1}^{J}A_{k}\bigg{)}\bm{u}_{s}^{1}}_{\bm{w}_{3}}.$ (67)
Now (54a) follows from (65), (54b) from (66), and (54c) from (67). $\blacksquare$

### v2

@@ I Introduction
  • Mathematical formulations (Sec. III-IV) for the stochastic learning of computational resource usage as graph-structured MSBPs where the graph structures arise from single or multi-core computational platforms. These formulations are motivated by maximum likelihood interpretations in the path space.
  • Numerical algorithms (Sec. V) for solving the proposed graph-stuctured MSBPs. We show that unlike generic graph structures, the MSBPs induced by single or multi-core computational resource usage lead to structured tensor optimization problems that can be leveraged to reduce the learning algorithms’ computational complexity from exponential to linear in number of observations.
  • Numerical experiments (Sec. VI) to illustrate the proposed learning framework. This includes a single-core MPC benchmark, and a synthetic multi-core software benchmark.
@@ V Algorithms
have guaranteed linear rate of convergenceThe proof sketch for discrete state space is as follows. The MSBP (34) can be recast [45, Sec. 4.1] as a Kullback-Leibler projection to a convex set that is an intersection of $|\Lambda|$ hyperplanes given by (34b). That this iterative Kullback-Leibler projection has guaranteed convergence follows from the seminal result on iterative Bregman projection [48], and from the fact that the Kullback-Leibler divergence is an instance of Bregman divergence. That the rate is linear follows from the co-ordinate descent analysis by Luo and Tseng [49, p. 25-26]. For the continuous state space, ref. [46, Thm. 4.7] proves the convergence of multi-marginal recursions. Ref. [47, Sec. 3] extends the result of [46] by showing linear rate of convergence., and the minimizer $\bm{M}^{\rm{opt}}$ for (34) is given in terms of the converged $\bm{U}$ as
@@ V-C Projections for Series-Parallel Graph-structured MSBP
$A_{k}:=K^{k,1}\left(\prod_{\sigma=2}^{s-1}\mathrm{diag}(\bm{u}_{\sigma}^{k})K^{k,\sigma}\right),\>B_{j}:=\bigodot_{k\neq j}A_{k}\>,$ (52)
and
$X_{\sigma}^{j}$ $:=K^{j,1}\bigg{(}\prod_{m=2}^{\sigma-1}\mathrm{diag}(\bm{u}_{m}^{j})K^{j,m}\bigg{)}\>,$
$Y_{\sigma}^{j}$ $:=K^{j,\sigma}\bigg{(}\prod_{m=\sigma+1}^{s-1}\mathrm{diag}(\bm{u}_{m}^{j})K^{j,m}\bigg{)}\>,$
$Z_{\sigma_{1},\sigma_{2}}^{j}$ $:=K^{j,\sigma_{1}}\bigg{(}\prod_{m=\sigma_{1}+1}^{\sigma_{2}-1}\mathrm{diag}(\bm{u}_{m}^{j})K^{j,m}\bigg{)}\>,$
(so $A_{k}=X_{s}^{k}=Y_{1}^{k}$), the projection (27) takes the form
${\rm{proj}}_{(1,1)}\!\left(\bm{K}\odot\bm{U}\right)\!$ $=\bm{u}_{1}^{1}\odot\left(\bigodot_{k=1}^{J}A_{k}\right)\bm{u}_{s}^{1},$ (53a)
${\rm{proj}}_{(1,s)}\!\left(\bm{K}\odot\bm{U}\right)\!$ $=\left({\bm{u}_{1}^{1}}^{\top}\cdot\bigodot_{k=1}^{J}A_{k}\right)^{\top}\odot\bm{u}_{s}^{1},$ (53b)
${\rm{proj}}_{(j,\sigma)}\!\left(\bm{K}\odot\bm{U}\right)\!$ $=\bm{u}_{\sigma}^{j}\!\odot\!\mathrm{diag}\Big{(}\!Y_{\sigma}^{j}\mathrm{diag}(\bm{u}_{s}^{1})B_{j}^{\top}\mathrm{diag}({\bm{u}_{1}^{1}})X_{\sigma}^{j}\!\Big{)},$ (53c)
@@
${\rm{proj}}_{(1,1),(j,2)}\!\left(\bm{K}\odot\bm{U}\right)\!=\mathrm{diag}\left(\bm{u}_{1}^{1}\right)$
$\quad\quad\cdot\Bigg{(}K^{j,1}\odot\bigg{(}B_{j}\mathrm{diag}(\bm{u}_{s}^{1}){Y_{2}^{j}}^{\top}\bigg{)}\Bigg{)}\cdot\mathrm{diag}\left(\bm{u}_{2}^{j}\right),$ (54a)
${\rm{proj}}_{(j,s-1),(1,s)}\!\left(\bm{K}\odot\bm{U}\right)\!=\mathrm{diag}\left(\bm{u}_{s-1}^{j}\right)$
$\quad\cdot\Bigg{(}K^{j,s-1}\odot\bigg{(}{X_{s-1}^{j}}^{\top}\mathrm{diag}(\bm{u}_{1}^{1})B_{j}\bigg{)}\Bigg{)}\cdot\mathrm{diag}\left(\bm{u}_{s}^{1}\right),$ (54b)
${\rm{proj}}_{(j,\sigma),(j,\sigma+1)}\!\left(\bm{K}\odot\bm{U}\right)\!=\mathrm{diag}\left(\!\bm{u}_{\sigma}^{j}\!\right)$
$\!\cdot\!\Bigg{(}K^{j,\sigma}\odot\bigg{(}{X_{\sigma}^{j}}^{\top}\mathrm{diag}(\bm{u}_{1}^{1})B_{j}\mathrm{diag}(\bm{u}_{s}^{1}){Y_{\sigma+1}^{j}}^{\!\!\!\!\!\!\top}\bigg{)}\Bigg{)}\!\cdot\!\mathrm{diag}\left(\!\bm{u}_{\sigma+1}^{j}\!\right).$ (54c)
Directly computing $A_{k}$ in (52) requires $1+2(s-2)$ matrix-matrix multiplications, each in this case involving $n^{3}$ floating-point operations. Thus, (53a) and (53b) each involve
$J(1+2(s-2))n^{3}+(J+1)n^{2}+n$ (55)
@@
$\left(J(1+2(s-2))+3\right)n^{3}+(J-1)n^{2}+n$ (56)
operations.
@@
In contrast, the expression (56)$=(2Js-3J+3)n^{3}+(J-1)n^{2}+n\!\geq\!n^{3}\!+\!n^{2}\!+\!n$. Thereby for $n$ large, Propositions 5 and 6 enable computation of unimarginal projections with $\mathcal{O}(Jsn^{2})$ complexity for the barycentric case, and $\mathcal{O}(Jsn^{3})$ for the series-parallel case. The $n^{3}$ complexity of the latter is a consequence of the treewidth ($=2$) of the series-parallel graph structure. Both cases have complexity linear in $Js$.
@@ VI-A2 Numerical Results
We computed (57) as the square roots of the optimal values of the corresponding Kantorovich linear programs [60, Ch. 3.1] that results from specializing (29) with $s=2$, $\varepsilon=0$. As expected, placing the marginals closer together in time results in higher accuracy in predictions.
@@ VI-B2 Numerical Results
$1$ $4.254\times 10^{-5}$ $1.020\times 10^{-7}$ $2.023\times 10^{-7}$ $1.412\times 10^{-7}$ $2.589\times 10^{-7}$
$2$ 00 $2.386\times 10^{-7}$ $2.329\times 10^{-7}$ $8.962\times 10^{-8}$ $1.908\times 10^{-7}$
$3$ 00 $2.392\times 10^{-7}$ $1.513\times 10^{-7}$ $4.693\times 10^{-8}$ $1.100\times 10^{-7}$
$4$ 00 $4.868\times 10^{-8}$ $2.050\times 10^{-7}$ $1.617\times 10^{-7}$ $1.204\times 10^{-7}$
@@ VII Conclusions
The learning-scheduling co-design that builds on the proposed method, will comprise our future work. Another direction of interest is to account for the asynchronous nature of the profiles by formulating the corresponding MSBPs with joint space-time stochasticity, i.e., as distributional generalization of optimal stopping problems initiated in the recent work [62]. Our multi-marginal formulation also opens up the possibility to incorporate additional flow rate constraints as in [63].
@@
$={\bm{u}_{1}^{1}}^{\top}\bigodot_{k=1}^{J}\Bigg{(}\underbrace{K^{k,1}\left(\prod_{\sigma=2}^{s-1}\mathrm{diag}(\bm{u}_{\sigma}^{k})K^{k,\sigma}\right)}_{:=A_{k}}\Bigg{)}\bm{u}_{s}^{1}$ (63)
$={\bm{u}_{1}^{1}}^{\top}\Bigg{(}\bigg{(}\underbrace{K^{j,1}\bigg{(}\prod_{m=2}^{\sigma-1}\mathrm{diag}(\bm{u}_{m}^{j})K^{j,m}\bigg{)}}_{:=X_{\sigma}^{j}}$
$\quad\cdot\mathrm{diag}(\bm{u}_{\sigma}^{j})\!\underbrace{K^{j,\sigma}\bigg{(}\!\prod_{m=\sigma+1}^{s-1}\!\!\!\!\!\mathrm{diag}(\!\bm{u}_{m}^{j}\!)\!K^{j,m}\!\!\bigg{)}}_{:=Y_{\sigma}^{j}}\!\!\bigg{)}\!\odot\!\underbrace{\bigodot_{k\neq j}A_{k}}_{:=B_{j}}\!\Bigg{)}\!\bm{u}_{s}^{1}$
$={\mathrm{trace}}\Big{(}\mathrm{diag}({\bm{u}_{1}^{1}})\bigr{(}X_{\sigma}^{j}\mathrm{diag}(\bm{u}_{\sigma}^{j})Y_{\sigma}^{j}\bigr{)}\mathrm{diag}(\bm{u}_{s}^{1})B_{j}^{\top}\Big{)}$
$={\bm{1}}^{\top}\mathrm{diag}\Big{(}\mathrm{diag}(\bm{u}_{\sigma}^{j})Y_{\sigma}^{j}\mathrm{diag}(\bm{u}_{s}^{1})B_{j}^{\top}\mathrm{diag}({\bm{u}_{1}^{1}})X_{\sigma}^{j}\Big{)}$
$=\underbrace{\bm{1}^{\top}}_{\bm{w}_{1}^{\top}}\cdot\mathrm{diag}(\bm{u}_{\sigma}^{j})\cdot\underbrace{\mathrm{diag}\Big{(}Y_{\sigma}^{j}\mathrm{diag}(\bm{u}_{s}^{1})B_{j}^{\top}\mathrm{diag}({\bm{u}_{1}^{1}})X_{\sigma}^{j}\Big{)}}_{\bm{w}_{2}},$ (64)
@@
$\langle\mathbf{K},\mathbf{U}\rangle={\rm{trace}}\Big{(}\mathrm{diag}(\bm{u}_{1}^{1})X_{\sigma_{1}}^{j}\mathrm{diag}(\bm{u}_{\sigma_{1}}^{j})$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\cdot Z_{\sigma_{1},\sigma_{2}}^{j}\mathrm{diag}(\bm{u}_{\sigma_{2}}^{j})Y_{\sigma_{2}}^{j}\mathrm{diag}(\bm{u}_{s}^{1})B_{j}^{\top}\Big{)}$
$=\bm{w}_{1}^{\top}\mathrm{diag}(\bm{u}_{1}^{1})\underbrace{\Bigg{(}K^{j,1}\odot\bigg{(}B_{j}\mathrm{diag}(\bm{u}_{s}^{1}){Y_{2}^{j}}^{\top}\bigg{)}\Bigg{)}}_{\Phi}\mathrm{diag}(\bm{u}_{2}^{j})\bm{w}_{3}$ (65)
$=\bm{w}_{1}^{\top}\mathrm{diag}(\bm{u}_{s-1}^{j})\!\underbrace{\Bigg{(}\!\!K^{j,s-1}\!\odot\!\bigg{(}\!\!{X_{s-1}^{j}}^{\!\!\!\!\!\!\top}\mathrm{diag}(\bm{u}_{1}^{1})B_{j}\!\!\bigg{)}\!\!\Bigg{)}}_{\Phi}\!\mathrm{diag}(\bm{u}_{s}^{1})\bm{w}_{3}$ (66)
$=\bm{w}_{1}^{\top}\mathrm{diag}(\bm{u}_{\sigma_{1}}^{j})\underbrace{\Bigg{(}\!\!Z_{\sigma_{1},\sigma_{2}}^{j}\!\odot\!\bigg{(}\!\!{X_{\sigma_{1}}^{j}}^{\!\!\top}\!\mathrm{diag}(\bm{u}_{1}^{1})B_{j}\mathrm{diag}(\bm{u}_{s}^{1}){Y_{\sigma_{2}}^{j}}^{\!\!\top}\!\bigg{)}\!\!\Bigg{)}}_{\Phi}$
$\quad\quad\quad\quad\quad\quad\cdot\mathrm{diag}(\bm{u}_{\sigma_{2}}^{j})\bm{w}_{3},$ (67)
where $\bm{w}_{1}=\bm{w}_{3}=\bm{1}$. Now (54a) follows from (65), (54b) from (66), and (54c) from (67). $\blacksquare$